Henry invested $12,000 in a new restaurant. When the restaurant was sold 2 years later, he received $27,000. Find his average annual returning by solving the equation 12,000(1+r)^2=27,000.
Answer by checkley77(12844)

12,000(1+r)^2=27,000
12,000(1+2r+r^2)=27,000
12,000+24,000r+12,000r=27,000
12,000r^2+24,000r+12,000-27,000=0
12,000r^2+24,000r-15,000=0
3,000(4r^2+8r-5)=0
3,000(2r+5)(2r-1)=0
2r-1=0
2r=1
r=1/2
r=.5 or 50% is the average annual return.
Proof:
12,000(1+.5)^2=27,000
12,000(1.5)^2=27,000
12,000*2.25=27,000
27,000=27,000